# Going beyond the creative limits of next-token prediction

### v2

@@ F.1 Ablation studies
Figure 21: Learning rates and algorithmic creativity for the SEDD (90M)model vs. GPT-2 (86M): MTP achieves higher algorithmic creativity than NTP when both are trained at their optimal learning rates.
@@
Figure 22: Task complexity and algorithmic creativity of SEDD (90M) model vs. GPT-2 (86M): MTP consistently outperforms NTP under varying task configurations, with some exceptions in the Line Constructionand Circle Constructiondatasets.
@@ F.2 Effect of seed length
Figure 23: GPT-2 (86M) Transformer achieves higher algorithmic creativity with longer seeds (top): We report algorithmic creativity with seeds of length 4 and 10, with both NTP and teacherless MTP. we do not see gains of seed-conditioning when it comes to diffusion training (right): We report algorithmic creativity of the SEDD (90M) model with or without seed-conditioning.
@@ F.3 Seed-conditioning for diffusion training
F.4 Format sensitivity for Triangle Discovery
@@
Figure 24: Sensitivity to formatting of the sequence in Triangle Discovery: We find that all our small models perform equally poorly with a node-wise representation of the input sequence, whereas there was a stark difference in performance with the edge-wise representation.
@@ Appendix G Additional experiments with medium-sized Transformer and SEDD
Figure 25: On a medium-sized ($\sim$400M) model, multi-token diffusion training improves algorithmic creativity from Eq 1 (top) on our four open-ended algorithmic tasks.
@@ Appendix H Decomposing creativity
Through following experiments on the GPT-2 (86M) model in the Sibling Discovery task, we try to understand the dynamics between two important quantities that affect algorithmic creativity: diversity/duplication and originality/memorization
@@ H.1 Diversity score
Figure 26: Algorithmic creativity and diversity are not necessarily correlated, exhibiting distinct dynamics: We find that NTP has a high diversity score through training, even higher than MTP. However, its algorithmic creativity reaches only a mediocre peak before descending, when MTP starts surpassing it.
@@ H.2 Decomposing algorithmic creativity as diversity and memorization
Better creativity can be achieved either by enhancing diversity or by reducing memorization – we try to disentangle these factors in this section. In Fig 27, we plot the algorithmic creativity, diversity, and memorization scores at the checkpoint of best algorithmic creativity. We see that seed-conditioning contributes to higher diversity but does little to bring down memorization; however, teacherless training contributes to higher diversity and also to reducing memorization. In Fig 26, we see that the best creativity and best diversity are not achieved at the same checkpoint.
@@ I.2 Effect of seed-conditioning
While these methods employ diverse strategies, a common feature is their reliance on multi-token objectives that capture broader dependencies across entire sequences. Representative examples include teacherless training (Bachmann & Nagarajan, 2024; Monea et al., 2023; Tschannen et al., 2023) and independent output heads or modules (Gloeckle et al., 2024; DeepSeek-AI et al., 2024) or inserting a lookahead attention (Du et al., 2023). Another line of research is discrete diffusion models (Hoogeboom et al., 2021; Austin et al., 2021; Gong et al., 2023; Lou et al., 2023), which avoid strict left-to-right factorization by iteratively refining an entire sequence at multiple positions. There are other models as well, such as energy-based models (Dawid & LeCun, 2023) and non-autoregressive models or (Gu et al., 2018).

### v3

@@ F.1 Ablation studies
Figure 21: Learning rates and algorithmic creativity for the SEDD (90M)model vs. GPT-2 (86M): MTP achieves higher algorithmic creativity than NTP when both are trained at their optimal learning rates. We use greedy decoding in this plot.
@@
Figure 22: Task complexity and algorithmic creativity of SEDD (90M) model vs. GPT-2 (86M): MTP consistently outperforms NTP under varying task configurations, with some exceptions in the Line Constructionand Circle Constructiondatasets. We use greedy decoding in this plot.
@@ F.2 Effect of seed length
Figure 23: GPT-2 (86M) Transformer achieves higher algorithmic creativity with longer seeds (top): We report algorithmic creativity with seeds of length 4 and 10, with both NTP and teacherless MTP (with greedy decoding). We do not see gains of seed-conditioning when it comes to diffusion training (right): We report algorithmic creativity of the SEDD (90M) model with or without seed-conditioning.
@@ F.3 Seed-conditioning for diffusion training
F.4 Effect of top-K sampling
Figure 24 reports the algorithmic creativity scores of NTP vs teacherless MTP training with seed-conditioning and top-$K$ sampling where K=50.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 24: Creativity scores for NTP vs teacherless MTP training with seed-conditioning + top-$K$ sampling where $K=50$.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 25: Memorization scores for NTP vs teacherless MTP training with seed-conditioning + top-$K$ sampling where $K=50$.
F.5 Format sensitivity for Triangle Discovery
@@
Figure 26: Sensitivity to formatting of the sequence in Triangle Discovery: We find that all our small models perform equally poorly with a node-wise representation of the input sequence, whereas there was a stark difference in performance with the edge-wise representation. We use greedy decoding in this plot.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 27: Seed conditioning improves creativity. “Best with seed” and “best without seed” stand for the best creativity score in each setting after we tune the sampling temperature from $\{0,0.5,1,2\}$.
F.6 Additional ablations
Figure 28 reports the ablations on GPT-2 (86M) for (1) NTP vs MTP, (2) effect of seed conditioning, (3) effect of top-$K$, and (4) effect of temperature sampling.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 28: Ablations on GPT-2 (86M) for (1) NTP vs teacherless MTP training, (2) effect of seed conditioning, (3) effect of top-K, and (4) effect of temperature sampling
@@ Appendix G Additional experiments with medium-sized Transformer and SEDD
Figure 29: On a medium-sized ($\sim$400M) model, multi-token diffusion training improves algorithmic creativity from Eq 1 (top) on our four open-ended algorithmic tasks. We use greedy decoding in this plot.
@@ Appendix H Decomposing creativity
Through following experiments on the GPT-2 (86M) model in the Sibling Discovery task, we try to understand the dynamics between two important quantities that affect algorithmic creativity: diversity/duplication and originality/memorization.
@@ H.1 Diversity score
Figure 30: Algorithmic creativity and diversity are not necessarily correlated, exhibiting distinct dynamics: We find that NTP has a high diversity score through training, even higher than MTP. However, its algorithmic creativity reaches only a mediocre peak before descending, when MTP starts surpassing it. We use greedy decoding in this plot.
@@ H.2 Decomposing algorithmic creativity as diversity and memorization
Better creativity can be achieved either by enhancing diversity or by reducing memorization – we try to disentangle these factors in this section. In Fig 31, we plot the algorithmic creativity, diversity, and memorization scores at the checkpoint of best algorithmic creativity. We see that seed-conditioning contributes to higher diversity but does little to bring down memorization; however, teacherless MTP training contributes to higher diversity and also to reducing memorization. In Fig 30, we see that the best creativity and best diversity are not achieved at the same checkpoint.
@@ I.2 Effect of seed-conditioning
While multi-token methods employ diverse strategies, a common feature is their reliance on objectives that capture dependencies across entire sequences. Representative examples include teacherless training (Bachmann & Nagarajan, 2024; Monea et al., 2023; Tschannen et al., 2023) and independent output heads or modules (Gloeckle et al., 2024; DeepSeek-AI et al., 2024) or inserting a lookahead attention (Du et al., 2023). Another line of research is discrete diffusion models (Hoogeboom et al., 2021; Austin et al., 2021; Gong et al., 2023; Lou et al., 2023), which avoid strict left-to-right factorization by iteratively refining an entire sequence at multiple positions. There are other models as well, such as energy-based models (Dawid & LeCun, 2023) and non-autoregressive models or (Gu et al., 2018). Frydenlund (2024) report gains from non-autoregressive models on the path-star task (Bachmann & Nagarajan, 2024), while Hu et al. (2025) propose a novel method call the Belief State Transformer that shows gains on these tasks. While we only report the effect of teacherless training on our tasks due to its simplicity, we speculate that these methods would show similar (or better) results as teacherless training in our settings.